how can i solve this problem????
the length of a rectangle is 2 cm more than four times the width if the perimeter of the rectangle is 24 cm what are its dimensions?

Answer by rapaljer(4671)

Let x = width of the rectangle
4x+2 = length of the rectangle

2 W + 2 L = P
2(x) + 2(4x+2) = 24
2x + 8x + 4 = 24
10x = 20
x= 2 cm = width
4x +2 =4(2) + 2 = 10 cm = length

Check:
2 W + 2 L = 24
2(2) + 2(10) = 24
4 + 20 = 24
It checks!!

R^2 at SCC